OpenStudy (book):

What is the value of the binary number 101 in base 10? 4 5 6 7

OpenStudy (steve816):

What class is this even for? These are some very random questions!

OpenStudy (steve816):

Reading from right to left in binary, the first digit is 1s, the second is 2s, the third is 4s, then 8s, 16s, etc. So 101 is (1*4)+(0*2)+(1*1) = 4+1 = 5.

OpenStudy (blazeryder):

(1*4)+(0*2)+(1*1) = 4+1 = 5

OpenStudy (book):

ty